# AgentFly: Fine-tuning LLM Agents without Fine-tuning LLMs

### v1

@@ 5.5 Ablation Studies
We analyse AgentFly’s hyper-parameter selection, component-wise Analysis, learning curves for both parametric and non-parametric case-based reasoning, out-of-distribution performance, and token costs.
@@
Table 3: The performance of AgentFly on the DeepResearcher dataset across different numbers of cases. We use gpt-$4.1$ as the planner and o44-mini as the executor.
@@ 5.5.1 Hyper-parameter Selection
AgentFly w/ Non-Parametric CBR 79.84 81.87 83.09 84.03 84.85
AgentFly w/ Parametric CBR 80.46 82.84 84.10 84.85 85.44
Table 4: Performance improvement of AgentFly over five learning iterations on the DeepResearcher dataset, demonstrating the benefit of accumulating cases in the Case Bank.
@@ 5.5.2 Component-wise Analysis
From Table 5, we observe a consistent pattern across HLE, SimpleQA, and DeepResearcher. Moving from an offline executor to live tools generally reduces hallucination and increases both F11 and PM, though the magnitude depends on task type (SimpleQA: +28.8+28.8 F11 / +63.3+63.3 PM, HLE: +4.8+4.8 / +7.1+7.1), and may even hurt on open-domain data (DeepResearcher: 18.0-18.0 / 2.1-2.1). Introducing planning (AgentFly w/o CBR) yields robust gains on each benchmark (HLE: +11.0+11.0 / +1.6+1.6, SimpleQA: +32.5+32.5 / +4.9+4.9, DeepResearcher: +29.1+29.1 / +11.5+11.5), indicating that explicit decomposition and tool orchestration systematically improve execution. Finally, case-based reasoning provides consistent, additive improvements (HLE: +4.5+4.5 / +7.0+7.0, SimpleQA: +3.7+3.7 / +5.3+5.3, DeepResearcher: +6.7+6.7 / +8.2+8.2). For HLE, however, without sufficient domain knowledge encoded in the backbone model, neither tool usage nor planning alone can reliably produce correct answers on long-tail, expert-level tasks. For DeepResearcher, we also identify data contamination (Shumailov et al., 2024) across the seven evaluated benchmarks, evidenced by a noticeable drop in both F11 and PM when moving from the offline executor to the online executor without planning (DeepResearcher: 18.0-18.0 F11 / 2.1-2.1 PM). This aligns with broader findings in the field (Sun et al., 2022, Yu et al., 2022, Zhou et al., 2025): simply using external knowledge can sometimes negatively affect the model, while the internal knowledge within the model plays an important role in QA tasks and can even outperform RAG.
@@ 5.5.3 Continual Learning Ability Boosted by Parametric and Non-Parametric CBR
Figure  1(c) and Table 4 present the continual learning curves across different memory designs for the AgentFly framework, comparing the performance of three configurations: AgentFly with non-parametric CBR or parametric CBR and AgentFly without CBR. The results demonstrate that the full AgentFly architecture consistently outperforms the ablated versions across all iterations, achieving higher accuracy at each step. Notably, removing CBR leads to a noticeable decline in performance, highlighting the effectiveness and complementary benefits of both parametric CBR and non-parametric CBR components in enhancing the continual learning capability of AgentFly. More importantly, we observe a learning curve of the accuracy on the DeepResearcher dataset with increased iterations, suggesting that memory-based approaches can effectively enhance LLM agents without requiring parameter updates.
@@ 5.5.4 Generalisation across Tasks
To assess out-of-distribution (OOD) generalisation, we follow the evaluation protocol of Zheng et al. (2025). Specifically, MusiQue (Trivedi et al., 2022), Bamboogle (Press et al., 2022), and PopQA (Mallen et al., 2022) are selected as OOD datasets due to their distinct question styles and information distributions, while NQ (Kwiatkowski et al., 2019), TQ (Joshi et al., 2017), HotpotQA (Yang et al., 2018), and 2Wiki (Ho et al., 2020) are used for training. We first collect and store trajectories from the training datasets in the case bank. During inference, AgentFly retrieves the four most relevant cases from the case bank for each target query. As shown in Figure 1(d), AgentFly achieves substantial improvements on all OOD benchmarks, with absolute gains ranging from 4.7% to 9.6%. These results highlight the effectiveness of case-based reasoning in enhancing generalisation to unseen tasks.
@@ 6 Discussion and Analysis
Building on the results in § 5 that establish the effectiveness of AgentFly, we further analyse its efficiency and operational behaviour. Specifically, we (i) analyse the average number of tool calls per task across three difficulty levels to assess how the MCP framework adapts as task complexity increases, (ii) characterise tool-call statistics, and (iii) evaluate the impact of using reasoning-oriented versus general-purpose models.
@@ 6.3 The Impact of Fast and Slow Think Mode
Table 6: The impact of fast and slow think mode on GAIA valiadation dataset (pass@1).
@@ 7 Conclusion
We introduce AgentFly, a memory-based learning paradigm that enables LLM agents to adapt online search without updating model weights. AgentFly formalises deep research agents as a memory-based Markov Decision Process (MDP) and implements it within a planner–executor framework, leveraging an episodic case bank to record and retrieve trajectories for continual policy improvement. Empirically, we achieve strong performance across GAIA, DeepResearcher, and SimpleQA. Ablation studies reveal that both parametric and non-parametric CBR are critical to the significant performance gains, and that a small, curated memory yields optimal results. These findings motivate future work on deep research tasks using memory-based MDP.

### v2

@@ 5.5 Ablation Studies
We analyse Memento’s hyper-parameter selection, component-wise Analysis, learning curves for both parametric and non-parametric case-based reasoning, out-of-distribution performance, and token costs.
@@
Table 3: The performance of Memento on the DeepResearcher dataset across different numbers of cases. We use gpt-$4.1$ as the planner and o44-mini as the executor.
@@ 5.5.1 Hyper-parameter Selection
Memento w/ Non-Parametric CBR 79.84 81.87 83.09 84.03 84.85
Memento w/ Parametric CBR 80.46 82.84 84.10 84.85 85.44
Table 4: Performance improvement of Memento over five learning iterations on the DeepResearcher dataset, demonstrating the benefit of accumulating cases in the Case Bank.
@@ 5.5.2 Component-wise Analysis
From Table 5, we observe a consistent pattern across HLE, SimpleQA, and DeepResearcher. Moving from an offline executor to live tools generally reduces hallucination and increases both F11 and PM, though the magnitude depends on task type (SimpleQA: +28.8+28.8 F11 / +63.3+63.3 PM, HLE: +4.8+4.8 / +7.1+7.1), and may even hurt on open-domain data (DeepResearcher: 18.0-18.0 / 2.1-2.1). Introducing planning (Memento w/o CBR) yields robust gains on each benchmark (HLE: +11.0+11.0 / +1.6+1.6, SimpleQA: +32.5+32.5 / +4.9+4.9, DeepResearcher: +29.1+29.1 / +11.5+11.5), indicating that explicit decomposition and tool orchestration systematically improve execution. Finally, case-based reasoning provides consistent, additive improvements (HLE: +4.5+4.5 / +7.0+7.0, SimpleQA: +3.7+3.7 / +5.3+5.3, DeepResearcher: +6.7+6.7 / +8.2+8.2). For HLE, however, without sufficient domain knowledge encoded in the backbone model, neither tool usage nor planning alone can reliably produce correct answers on long-tail, expert-level tasks. For DeepResearcher, we also identify data contamination (Shumailov et al., 2024) across the seven evaluated benchmarks, evidenced by a noticeable drop in both F11 and PM when moving from the offline executor to the online executor without planning (DeepResearcher: 18.0-18.0 F11 / 2.1-2.1 PM). This aligns with broader findings in the field (Sun et al., 2022, Yu et al., 2022, Zhou et al., 2025): simply using external knowledge can sometimes negatively affect the model, while the internal knowledge within the model plays an important role in QA tasks and can even outperform RAG.
@@ 5.5.3 Continual Learning Ability Boosted by Parametric and Non-Parametric CBR
Figure  1(c) and Table 4 present the continual learning curves across different memory designs for the Memento framework, comparing the performance of three configurations: Memento with non-parametric CBR or parametric CBR and Memento without CBR. The results demonstrate that the full Memento architecture consistently outperforms the ablated versions across all iterations, achieving higher accuracy at each step. Notably, removing CBR leads to a noticeable decline in performance, highlighting the effectiveness and complementary benefits of both parametric CBR and non-parametric CBR components in enhancing the continual learning capability of Memento. More importantly, we observe a learning curve of the accuracy on the DeepResearcher dataset with increased iterations, suggesting that memory-based approaches can effectively enhance LLM agents without requiring parameter updates.
@@ 5.5.4 Generalisation across Tasks
To assess out-of-distribution (OOD) generalisation, we follow the evaluation protocol of Zheng et al. (2025). Specifically, MusiQue (Trivedi et al., 2022), Bamboogle (Press et al., 2022), and PopQA (Mallen et al., 2022) are selected as OOD datasets due to their distinct question styles and information distributions, while NQ (Kwiatkowski et al., 2019), TQ (Joshi et al., 2017), HotpotQA (Yang et al., 2018), and 2Wiki (Ho et al., 2020) are used for training. We first collect and store trajectories from the training datasets in the case bank. During inference, Memento retrieves the four most relevant cases from the case bank for each target query. As shown in Figure 1(d), Memento achieves substantial improvements on all OOD benchmarks, with absolute gains ranging from 4.7% to 9.6%. These results highlight the effectiveness of case-based reasoning in enhancing generalisation to unseen tasks.
@@ 6 Discussion and Analysis
Building on the results in § 5 that establish the effectiveness of Memento, we further analyse its efficiency and operational behaviour. Specifically, we (i) analyse the average number of tool calls per task across three difficulty levels to assess how the MCP framework adapts as task complexity increases, (ii) characterise tool-call statistics, and (iii) evaluate the impact of using reasoning-oriented versus general-purpose models.
@@ 6.3 The Impact of Fast and Slow Think Mode
Table 6: The impact of fast and slow think mode on GAIA validation dataset (pass@1).
@@ 7 Conclusion
We introduce Memento, a memory-based learning paradigm that enables LLM agents to adapt online search without updating model weights. Memento formalises deep research agents as a memory-based Markov Decision Process (MDP) and implements it within a planner–executor framework, leveraging an episodic case bank to record and retrieve trajectories for continual policy improvement. Empirically, we achieve strong performance across GAIA, DeepResearcher, and SimpleQA. Ablation studies reveal that both parametric and non-parametric CBR are critical to the significant performance gains, and that a small, curated memory yields optimal results. These findings motivate future work on deep research tasks using memory-based MDP.